OpenStudy (anonymous):

ln(3) -2ln(4+8) use the properties of log to write single log with a coefficient of 1....please show work

OpenStudy (anonymous):

2ln(4+8) = ln 12^2 ln3 - ln 12^2 = ln 3/144 = ln 1/48

OpenStudy (anonymous):

ty very much

OpenStudy (anonymous):

yw